OpenStudy (anonymous):

6 coins are tossed simultaneously whats the probability of getting at least 6 heads?

OpenStudy (anonymous):

HHHHHH -> \(\huge \frac{1}{2^6} \)

OpenStudy (anonymous):

at least?

OpenStudy (lgbasallote):

there could be more :p

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

more than 6 heads on six coins? hmmm

OpenStudy (anonymous):

fool for math

OpenStudy (anonymous):

can someone help me on my question please? it's really hard

OpenStudy (lgbasallote):

i see no reason why not

Parth (parthkohli):

You multiply the probability of each event when you have questions like these where the word 'and' or 'simultaneous' is used.

OpenStudy (anonymous):

unless there is a typo, at least 6 heads means exactly 6 heads and foolformath has it

OpenStudy (anonymous):

(1/2)^6

Parth (parthkohli):

At least means 6..

OpenStudy (cwrw238):

hmm - double headed coins!

Parth (parthkohli):

How can there be more than 6?

OpenStudy (lgbasallote):

those are hydra coins

OpenStudy (cwrw238):

lol

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow ({1 \over 2})^{6} }\)

OpenStudy (anonymous):

can someone help me with my normal distribution problem please?

Parth (parthkohli):

Magical coins lol

OpenStudy (anonymous):

it's really hard

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

probability of head on any one tossis \(\frac{1}{2}\) probability of getting 6 in a row is \((\frac{1}{2})^6=\frac{1}{2^6}=\frac{1}{64}\)

OpenStudy (anonymous):

satellite, can you help me out please?

OpenStudy (anonymous):

i really need your help

OpenStudy (anonymous):

whats the question?

Parth (parthkohli):

Mr. Satellite is our expert. Don't underestimate him ;P

OpenStudy (anonymous):

i'll bump it in a few secs

OpenStudy (anonymous):

bumped

OpenStudy (anonymous):

probability of getting atleast 6 heads = (number of possible ways of getting number of heads greater than or equal to 6) /(total possibilities that can occur in the given process) = (1)/(2x2x2x2x2x2) = \[2^{-6}\]